OpenStudy (anonymous):

Using the graph below, calculate the average rate of change for f(x) from x = 2 to x = 6. exponential function going through points 2, 2 and 6, 6 x = −4 x = −1 x = 1 x = 4

OpenStudy (imstuck):

Do you have the y coordinates to go with those x coordinates?

OpenStudy (freckles):

you need evaluate \[\frac{f(6)-f(2)}{6-2}\]

OpenStudy (anonymous):

I dont have the Y coordinate to go with the x's

OpenStudy (anonymous):

@freckles would it be 1 then?

OpenStudy (freckles):

Well you need the corresponding y -value for x=6 that is what is meant be the notation f(6) You also need the corresponding y-value for x=2 that is what is meant be the notation f(2) are you saying f(6) is 6 and f(2) is 2?

OpenStudy (anonymous):

yes, I think so if I understand what you're saying, I'm still slightly confused.

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@Hero ?

OpenStudy (anonymous):

@sammixboo ?

OpenStudy (freckles):

confused on what part?

OpenStudy (freckles):

Are you given in ordered pairs? ordered pairs (x,y) or you can write it as (x,f(x))

OpenStudy (freckles):

I'm sure what 2,2 means do you mean (2,2)?

OpenStudy (freckles):

(2,2) means (2,f(2)=2) that means when x=2, f(2)=2

OpenStudy (anonymous):

I did the first equation you put up and got one. I'll upload the full problem

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

I can't get that to load but if f(2)=2 and f(6)=6 then yes the average rate of change is 1 since (6-2)/(6-2)=4/4=1.